# African Data Ethics: A Discursive Framework for Black Decolonial Data Science

### v1

@@ 4.5. Data Self-Determination
African data science should be an avenue for bolstering the self-determination of Indigenous African communities.
@@
“For Africans, By African”. This principle is inspired by the concerted efforts of African data scientists to reclaim leadership in African data science work (Chan et al., 2021). To combat deficit-based narratives about Africa, African data scientists need to reclaim and celebrate their strength, rich cultures, and scientific achievements in conducting RDS (Abebe et al., 2021; Adelani et al., 2022; Hountondji, 2004; Coetzee, 2004b; Lauer, 2017; Gwagwa, 2019; Carman and Rosman, 2023). The diverse values and perspectives of African communities should ground the development of African data ethics (Coetzee and Roux, 2004; Segun, 2021; African Union, 2024; Ruttkamp-Bloem, 2023; Gwagwa et al., 2022; Dignum, 2023; Olojede, 2023). Given the thousands of cultures that comprise Africa, the potential for novel approaches to data science must be explored (Eke and Ogoh, 2022; Goffi, 2023; Coetzee, 2004a; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Day and Rennie, 2023; Kohnert, 2022).
@@
African data should not be primarily collected for Western tech powers or published for immediate and uncontrolled use (Birhane, 2020; Hountondji, 2004). African data practitioners do not need tech superpowers to speak for Africans on the global stage, provide pre-trained models, or ensure work meets the standards of Western data institutions, Africans are more than capable of leading without interference (Ndjungu, 2020; Abebe et al., 2021; Goffi, 2023; Mhlambi and Tiribelli, 2023; Okolo et al., 2023; Ade-Ibijola and Okonkwo, 2023; Biko, 2004). This does not mean Africans should not collaborate with external data practitioners and vice versa (Hountondji, 2004; Eke et al., 2023b). Rather, local African data practitioners must lead data science work so its development is properly situated in the communities it will be used (Nwankwo and Sonna, 2019; Lauer, 2017; Kiemde and Kora, 2022; Eke et al., 2023a).
@@
Treasure Indigenous Knowledge. DDS should preserve, center, and continue the development of Indigenous knowledge. With the legacy of colonial epistemic injustice, African modernization and Indigenous knowledge preservation are often viewed as at odds with each other (African Union, 2024; Kohnert, 2022; Eke et al., 2023b). On the contrary, many documents hold Indigenous knowledge as a pivotal component of RDS in Africa.
@@
DDS can be used to store Indigenous languages, customs, and history in close consultation with Indigenous communities. However, some are concerned that joining a globalized data ecosystem will lead to a loss of culture and identity (African Union, 2024; Abebe et al., 2021; Eke et al., 2023a; Ade-Ibijola and Okonkwo, 2023). As elders, griots, and other stewards of Indigenous knowledge pass, younger generations have to take on the responsibility of preserving their community’s culture (Kotut et al., 2024; Ramose, 2004b). There are over 1500 languages indigenous to Africa, but very few are represented in data technology, such as natural language processing (NLP), which leaves out large portions of Africans from using technology (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023). Pre-colonial Indigenous knowledge needs to be reclaimed to develop African data values that reflect local communities (Abdul, 2023; Chan et al., 2021). Local communities can never fully be represented if there is not an understanding of their roots or history (Ramose, 2004b). Building datasets that represent Indigenous languages for inclusive models opens a whole set of new users who can digitally store and analyze Indigenous knowledge that is typically shared orally for future generations (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Moahi, 2007). The boundaries on what Indigenous knowledge should be a part of DDS must be understood by consulting with the community before proceeding on any project (Kotut et al., 2024; Moahi, 2007).
@@
African philosophers emphasize Indigenous knowledge isn’t limited to the past (Hountondji, 2004). Investing in African RDS is an investment in creating new Indigenous knowledge (Uzomah et al., 2023; Lauer, 2017). Local talent does not have to reinvent the wheel to explore open questions in the more recent field of data science (Lauer, 2017; Mabe, 2007; Moahi, 2007). The richness of African knowledge can develop new RDS practices and understandings (Abdul, 2023; Mhlambi, 2020; Biko, 2004; Coetzee, 2004a).
@@
Data Sovereignty & Privacy. Mechanisms must be developed to protect African creativity and privacy in the development of DDS. Given the legacies of extractive colonialism, ownership is viewed as the key to data sovereignty in Africa (Gwagwa, 2019; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Kiemde and Kora, 2022). African ownership in the data science process can be achieved by codifying intellectual property rights (African Union, 2024), enforcing data ownership (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023), and exploring Indigenous conceptions of collective privacy (Nwankwo and Sonna, 2019; Goffi, 2023; Mabe, 2007; Langat et al., 2020; Moahi, 2007).
@@
Africans are often regarded as “simply” data subjects (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023). The role of a data subject is materially essential to data science work (without data, nothing can be done). The narratives of collecting as much data as possible to achieve generalizability devalue data subjects as dehumanized resources (Gwagwa, 2019; Birhane, 2020; Olojede, 2023; Mhlambi, 2020; Ndjungu, 2020). This devaluing encourages data collectors to share and use data without any knowledge, consent, or compensation of data subjects (Sinha and Zulfa, 2023; Nyerere, 1962). Many African data ethicists call for a correction of this narrative to recognize data subjects as the proper owners of their data by shifting power and access control to data subjects (Day and Rennie, 2023; Ruttkamp-Bloem, 2023; Abdul, 2023). Achieving this shift in ownership should be done by demanding data-sharing terms and not working with data collaborators who do not honor these terms (Biko, 2004; Okolo et al., 2023). African ownership of data, resources for data science, and technical contributions should be non-negotiable for RDS.
4.6. Invest in Data Institutions & Infrastructures.
Prioritizing infrastructure, investing in people, and establishing sound policy and governance frameworks should be measured to not deter social progress in the name of technological progress.
Measured Development. The development of data science ecosystems should be balanced, measured, inclusive, community-minded, and holistic (Kohnert, 2022; Eke et al., 2023a; Ade-Ibijola and Okonkwo, 2023). Without this approach, the adoption of AI and other data technologies can lead to more unrest and inequality across Africa. To many, the potential of DDS is profound and would change the trajectory of African development (African Union, 2024; Mabe, 2007). Data are viewed as the driving resource for the Fourth Industrial Revolution (Carman and Rosman, 2023; Gwagwa, 2019). There are African data scientists and governments who insist joining the AI boom will provide Africa the quality of life benefits afforded to the major players of past industrial revolutions (Okolo et al., 2023; Kohnert, 2022; Coetzee, 2004a). However, there is skepticism about wholeheartedly diving into large-scale data science adoption (Uzomah et al., 2023; Olojede, 2023). There is a need to quell the AI hype as the solution for all African problems and consider who will actually be served: Africans or the external powers propelling the AI boom (Wareham, 2021; Birhane, 2020; Sinha and Zulfa, 2023). Through the paradigm of measured development, technical development should move at the pace of social development (Kiemde and Kora, 2022; Nyerere, 1962). Paulin Hountondji’s critique of science in Africa applies well to data science development. Development should not be driven by “scientific extroversion” or catching up with the West (Hountondji, 2004; Goffi, 2023). Rather, the development of data science should be an investment in the progress of African people based on African intellect, priorities, and visions of the future (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Biko, 2004).
@@
Technical Infrastructure. To implement DDS in Africa, practitioners call for investment in physical data science infrastructure, assessment of the current capacities of technical infrastructure, and development of responsible data management practices (Moahi, 2007; Ruttkamp-Bloem, 2023). Achieving this principle in Africa is a big feat when electricity and broadband access is not only sparse but one of the most costly to access in the world (Okolo et al., 2023; Ade-Ibijola and Okonkwo, 2023). Nigeria, Mozambique, and Rwanda have recognized the need to invest in technical infrastructure and have partnered with external tech companies and international financial institutions to build their respective capacity to host DDS (Okolo et al., 2023). There are also innovative ways to work with current technical infrastructure to lessen reliance on external investment (Mhlambi, 2020). Technical infrastructure development should also coincide with the development of responsible data management protocols so that African data and data science work are not vulnerable to dispossession (Abebe et al., 2021; Gwagwa, 2019).
@@
Governance Infrastructure. We need sustainable and measured governance infrastructure to guide the development of DDS (African Union, 2024; Chan et al., 2021). Policy measures and regulations are major priorities for African data science communities to guide RDS practices. African Union member states are slowly developing data protection regulations, but many documents stress the urgency for African data policy (Kiemde and Kora, 2022; Plantinga et al., 2024). Without clear policies and legal standards for RDS, African data scientists lack guidance in their practices leaves African communities vulnerable to data exploitation from external and internal actors alike (Abebe et al., 2021; Cisse, 2018; Mandaza, 2004). Governance infrastructures include incremental regulations (Gwagwa, 2019), monitoring bodies (Goffi, 2023), continental commitments (African Union, 2024), and algorithmic impact assessments (Sinha and Zulfa, 2023).
Support Formal & Informal Collectives. Capacity-building in Africa necessitates the support of diverse data science collectives (Okolo et al., 2023; Abebe et al., 2021). 81% of jobs in Africa are based in informal economies (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023). As such, only focusing on supporting data science research (in which wider recognition and acceptance is a common issue related to epistemic injustice (Eke and Ogoh, 2022; Chan et al., 2021)) neglects a large portion of potential data collaborators (Hountondji, 2004). There should be efforts to connect Africans interested in using data science for entrepreneurship (Biko, 2004; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023) and accessible data science job training (Abebe et al., 2021). However, these collectives should not be siloed. The boundary between formal and informal data organizations should be dismantled to exchange technical knowledge, coordinate work, and pool resources (Dieng, 2023; Kling et al., 2023). Both forms of data collectives have important functions and need to rely on each other to flourish. One form of collective is not meant to replace the other (Osaghae, 2004). If both of these collectives are not supported, African data scientists will have to seek support outside of their communities, which furthers the “brain drain” of highly skilled Africans to the West (Okolo et al., 2023). Investing in collectives also builds a workforce for in-house development which reduces foreign dependence (Kiemde and Kora, 2022; Carman and Rosman, 2023; Plantinga et al., 2024).
4.7. Prioritize Education & Youth
Youth involvement and education are essential for ensuring the continued development and implementation of ethical data science practices that respect cultural contexts, African philosophy, and Indigenous knowledge.
@@ 4.7. Prioritize Education & Youth
Holistic Education. The African population has low attainment of digital skills (Okolo et al., 2023; Ade-Ibijola and Okonkwo, 2023). As large foreign technology companies set root in Africa, policymakers stress the need for monumental efforts to train local talent (African Union, 2024; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Mhlambi, 2020; Cisse, 2018). Providing technical skills early in education will help prepare a strong cohort of future data scientists (Sinha and Zulfa, 2023; Nwankwo and Sonna, 2019). There should also be investments in integrating AI curricula in informal organizations like the Data Values Project to reduce educational barriers (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023). Importantly, an indispensable part of a comprehensive data science education is data ethics (Goffi, 2023; Kiemde and Kora, 2022; Ramose, 2004b). An United Nations Educational, Scientific and Cultural Organization (UNESCO) survey found that very few African countries feel equipped to contend with the ethical implications of AI (Kiemde and Kora, 2022). Teaching data ethics in Africa should involve centering the lived experiences and culture of the students (Goffi, 2023; Kiemde and Kora, 2022). Students should be educated about the common dangers of data science and also develop their ethical discernment to prepare them for the sociotechnical complexities of data science.

### v3

@@ 4.2. Assert Data Self-Determination
Responsible African data science should be an avenue for bolstering the self-determination of Indigenous African communities.
@@
“For Africans, By Africans”. This principle is inspired by the concerted efforts of African data scientists to reclaim leadership in African data science work (Chan et al., 2021). To combat deficit-based narratives about Africa, scholars in our corpus call on African data scientists to reclaim and celebrate their strength, rich cultures, and scientific achievements in conducting RDS (Abebe et al., 2021; Adelani et al., 2022; Hountondji, 2004; Coetzee, 2004b; Lauer, 2017; Gwagwa, 2019; Carman and Rosman, 2023). Given the thousands of cultures that comprise Africa, African data scientists believe that grounding African data technology development in local talent, knowledge and data ethics can address local problems effectively and at scale (Eke and Ogoh, 2022; Goffi, 2023; Coetzee, 2004a; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Day and Rennie, 2023; Kohnert, 2022; Coetzee and Roux, 2004; Segun, 2021; African Union, 2024; Ruttkamp-Bloem, 2023; Gwagwa et al., 2022; Dignum, 2023; Olojede, 2023).
@@
To achieve the ideal of African-led data science, the authors in our corpus highlight several changes to how African data scientists approach their work. African data should not be primarily collected for technology powers or published for immediate and uncontrolled use (Birhane, 2020; Hountondji, 2004). Additionally, African data practitioners do not need tech superpowers to speak for Africans on the global stage, only provide off-the-shelf models, or oversee the standards of African data science institutions; Africans are more than capable of leading without interference (Ndjungu, 2020; Abebe et al., 2021; Goffi, 2023; Mhlambi and Tiribelli, 2023; Okolo et al., 2023; Ade-Ibijola and Okonkwo, 2023; Biko, 2004). This does not mean Africans should not collaborate with external data practitioners and vice versa (Hountondji, 2004; Eke et al., 2023b). Rather, local African data practitioners must lead data science work so its development is properly situated in the communities in which the technology will be deployed (Nwankwo and Sonna, 2019; Lauer, 2017; Kiemde and Kora, 2022; Eke et al., 2023a).
@@
Treasure Indigenous Knowledge. With the legacy of colonial epistemic injustice, African modernization and Indigenous knowledge preservation are often viewed as at odds with each other (African Union, 2024; Kohnert, 2022; Eke et al., 2023b). On the contrary, many documents hold Indigenous knowledge as a pivotal component of responsible data science in Africa. As elders, griots, and other stewards of Indigenous knowledge pass, younger generations are obligated to preserve their community’s culture (Kotut et al., 2024; Ramose, 2004b).
@@
Data-driven technology can be used to store Indigenous languages, customs, and history in close consultation with Indigenous communities. There are over 1500 languages indigenous to Africa, but very few are represented in data technology, such as natural language processing (NLP), which leaves out large portions of Africans from using ubiquitous technology (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023). Local communities can never fully be represented if there is not an understanding of their roots or history (Ramose, 2004b). Building datasets that represent Indigenous languages for inclusive models opens a whole set of new users who can digitally store and analyze Indigenous knowledge that is typically shared orally for future generations (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Moahi, 2007). However, some scholars caution that releasing Indigenous information into the globalized data ecosystem could lead to a loss of cultural control and appropriation (African Union, 2024; Abebe et al., 2021; Eke et al., 2023a; Ade-Ibijola and Okonkwo, 2023). The boundaries of what Indigenous knowledge should be a part of data-driven technology must be understood by consulting with the community before proceeding on any project (Kotut et al., 2024; Moahi, 2007).
@@
African philosophers also emphasize Indigenous knowledge isn’t limited to the past (Hountondji, 2004). Pre-colonial Indigenous knowledge needs to be reclaimed to develop African data values that reflect local communities (Abdul, 2023; Chan et al., 2021). Investing in African responsible data science is an investment in creating new Indigenous knowledge (Uzomah et al., 2023; Lauer, 2017). Local talent do not have to reinvent the wheel to explore open questions in the more recent field of data science (Lauer, 2017; Mabe, 2007; Moahi, 2007). The richness of African knowledge can develop new responsible data science practices and understandings (Abdul, 2023; Mhlambi, 2020; Biko, 2004; Coetzee, 2004a).
@@
Protect Data Agency. Given the legacies of extractive colonialism, ownership is viewed as the key to data self-determination in Africa (Gwagwa, 2019; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Kiemde and Kora, 2022). African ownership in the data science process can be achieved by codifying intellectual property rights (African Union, 2024), enforcing data ownership (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023), and exploring Indigenous conceptions of collective privacy (Nwankwo and Sonna, 2019; Goffi, 2023; Mabe, 2007; Langat et al., 2020; Moahi, 2007).
@@
Africans are often regarded as “simply” data subjects (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023). However, the role of a data subject is materially essential to data science work (without data, nothing can be done). Performance progress narratives that require unbounded data consumption devalue data subjects as dehumanized sources of raw material (Gwagwa, 2019; Birhane, 2020; Olojede, 2023; Mhlambi, 2020; Ndjungu, 2020). This devaluing encourages data collectors to share and use data without the knowledge, consent, or compensation of data subjects (Sinha and Zulfa, 2023; Nyerere, 1962). Many African data ethicists call for a correction of this narrative to recognize data subjects as the proper owners by shifting power and access control to data subjects (Day and Rennie, 2023; Ruttkamp-Bloem, 2023; Abdul, 2023). Achieving this shift in ownership should be done by demanding data-sharing terms and not working with data collaborators who do not honor these terms (Biko, 2004; Okolo et al., 2023). African ownership of data, resources for data science, and technical contributions are a non-negotiable for responsible data science in Africa.
4.3. Invest in Local Data Institutions & Infrastructures
Prioritizing infrastructure, investing in local talent, and establishing sound policy and governance frameworks are essential for sustained and independent RDS in Africa.
@@ 4.3. Invest in Local Data Institutions & Infrastructures
Invest in Physical & Organizational Infrastructures. To implement data technology in Africa, practitioners call for investment in physical data science infrastructure, assessment of the current capacities of technical infrastructure, and development of responsible data management practices (Moahi, 2007; Ruttkamp-Bloem, 2023). Achieving this principle in Africa is a big feat when electricity and broadband access is not only sparse but one of the most expensive to access in the world (Okolo et al., 2023; Ade-Ibijola and Okonkwo, 2023). Nigeria, Mozambique, and Rwanda have recognized the need to invest in technical infrastructure and have partnered with external tech companies and international financial institutions to build their respective capacities to host data-driven systems (Okolo et al., 2023). There are also innovative ways to work with current technical infrastructure to lessen reliance on external investment (Mhlambi, 2020). Technical infrastructure development should also coincide with the development of responsible data management protocols so African data and data science work are not vulnerable to dispossession (Abebe et al., 2021; Gwagwa, 2019).
@@
Cultivate Governance Infrastructure. The authors in our corpus call for sustainable and measured governance infrastructure to guide the development of data science in Africa (African Union, 2024; Chan et al., 2021). Policy measures and regulations are major priorities for African data science communities. African Union member states are slowly developing data protection regulations, but many documents stress the urgency for African data policy (Kiemde and Kora, 2022; Plantinga et al., 2024). Without clear policies and legal standards for RDS, African data scientists lack guidance in their practices, leaving African communities vulnerable to data exploitation from external and internal actors alike (Abebe et al., 2021; Cisse, 2018; Mandaza, 2004). Governance infrastructures include incremental regulations (Gwagwa, 2019), monitoring bodies (Goffi, 2023), continental commitments (African Union, 2024), and algorithmic impact assessments (Sinha and Zulfa, 2023).
@@
Support Formal & Informal Data Science Collectives. The African population has low attainment of digital skills (Okolo et al., 2023; Ade-Ibijola and Okonkwo, 2023). As large foreign technology companies set root in Africa, policymakers stress the need for monumental efforts to train local talent (African Union, 2024; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Mhlambi, 2020; Cisse, 2018). Providing technical skills early in education will help prepare a strong cohort of future data scientists (Sinha and Zulfa, 2023; Nwankwo and Sonna, 2019). An indispensable part of a comprehensive data science education is data ethics (Goffi, 2023; Kiemde and Kora, 2022; Ramose, 2004b). A United Nations Educational, Scientific and Cultural Organization (UNESCO) survey found that very few African countries feel equipped to contend with the ethical implications of AI (Kiemde and Kora, 2022). Teaching data ethics in Africa should involve centering the lived experiences and cultures of the students (Goffi, 2023; Kiemde and Kora, 2022). Students should be educated about the common harms of data science and also develop their ethical discernment to prepare them for the sociotechnical complexities of data science.
However, only focusing on supporting formal data science education (in which wider recognition and acceptance is a common issue related to epistemic injustice (Eke and Ogoh, 2022; Chan et al., 2021)) neglects a large portion of potential data collaborators (Hountondji, 2004). 81% of jobs in Africa are based in informal economies (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023). There should also be investments in integrating AI curricula in informal organizations like the Data Values Project to reduce educational barriers (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023). There should be efforts to connect Africans interested in using data science for entrepreneurship (Biko, 2004; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023) and accessible data science job training (Abebe et al., 2021). Capacity-building in Africa necessitates the support of formal and informal data science collectives (Okolo et al., 2023; Abebe et al., 2021).
Crucially, scholars in our corpus recommend dismantling the boundary between formal and informal data organizations to exchange technical knowledge, coordinate work, and pool resources (Dieng, 2023; Kling et al., 2023). Both forms of collectives have vital affordances and need to rely on each other to flourish. One form of collective is not meant to replace the other (Osaghae, 2004). If both of these collectives are not supported, African data scientists will have to seek support outside of their communities, which furthers the “brain drain” of highly skilled Africans to global superpowers (Okolo et al., 2023). Investing in collectives also builds a workforce for in-house development which reduces foreign dependence (Kiemde and Kora, 2022; Carman and Rosman, 2023; Plantinga et al., 2024).
4.4. Utilize Communalist Practices
The development and deployment of data-driven technology should mitigate harms, involve communities in decision-making, and ensure reciprocal benefits for African stakeholders.
Ground Design in Community-Engaged Consensus. Consensus-building is a well-practiced strategy from African communities that can inform responsible practices and encourage effective collaboration (Wiredu, 2004b). Rather than majority-rule common in European societies, African elders discuss issues until they all agree on a final decision (Wiredu, 2004a; Carman and Rosman, 2023). Achieving consensus requires the final decision to be 1) the dominant view of the group, 2) in line with the common good, and 3) aligned with the morals of the individual parties (Coetzee, 2004b). Consensus should be broached in an environment of trust, practical reason, humility, openness, and respect for the viewpoints of all involved parties (Coetzee, 2004b; Nwankwo and Sonna, 2019; Gwagwa, 2019; Mhlambi, 2020; Okolo et al., 2023; Gwagwa et al., 2022).
Community engagement provides spaces for consensus in the data science lifecycle to include more perspectives (Day and Rennie, 2023; Mabe, 2007). Akan philosophies regard the community as an invaluable resource that guides how every individual lives (Wiredu, 2004a; Metz, 2021; Coetzee, 2004b; Mhlambi and Tiribelli, 2023; Gwagwa et al., 2022). Therefore, community input is crucial for constructing a full picture of technical requirements, especially in high-stakes domains (Sinha and Zulfa, 2023; Mhlambi, 2020; Eke et al., 2023a). The concept of community can be misappropriated to deem any collection of stakeholders as sufficient community representatives. African communitarian ethics define a community as individuals with a shared identity who are emotionally invested in each other (Nwankwo and Sonna, 2019; Sinha and Zulfa, 2023; Ruttkamp-Bloem, 2023; Gyekye, 2004). With this more narrow definition of community, involving affected communities in all stages of the lifecycle requires building trust and respecting boundaries by gaining an understanding of cultural norms (Abebe et al., 2021; Ade-Ibijola and Okonkwo, 2023). Additionally, community members should be sufficiently trained or educated on the nature of the technology so they can provide well-informed input (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Adelani et al., 2022; Plantinga et al., 2024).
Consensus processes should also include procedures for documentation to keep track of disagreements, dissenting opinions, and the progression of project values (Kling et al., 2023). It is not easy to achieve these conditions, so conflict management, negotiation, and reasonable bargaining are helpful mechanisms to fully consider and resolve contradicting positions (Gwagwa, 2019; Osaghae, 2004; Dieng, 2023). Consensus should be a dynamic feedback loop to ensure every contributor is on the same page about the team’s approach to RDS (Segun, 2021; Nwankwo and Sonna, 2019; Uzomah et al., 2023; Kohnert, 2022; Abebe et al., 2021; Gwagwa, 2019; Dignum, 2023). Community-centered consensus-building is a co-creation process in which all stakeholders depend on each other (Langat et al., 2020; Kohnert, 2022; Abebe et al., 2021; Adelani et al., 2022; Nwankwo and Sonna, 2019; Lauer, 2017; Kiemde and Kora, 2022).
Instantiate Reciprocal Relationships. In many African philosophies, reciprocity is the foundation of a healthy society. In Akan society, practicing reciprocity ensures that community needs are met, while building deep social bonds (Wiredu, 2004b). African perfectionist proponents go as far as to assert that assisting others in achieving their goals makes someone more of a person (Wareham, 2021; Wiredu, 2004b). Without reciprocity, society becomes imbalanced and co-dependent (Coetzee, 2004b; Mhlambi and Tiribelli, 2023; Nyerere, 1962). There are numerous examples of African data subjects not reaping any benefits from the data collaborations they participate in (Gwagwa, 2019; Abebe et al., 2021). This often leads to technically mediated harms while the controllers of data amass profits (Gwagwa, 2019). Therefore, sustainable RDS should practice reciprocity on several dimensions (Wiredu, 2004b). If someone contributes to DDT they should meaningfully benefit from the system or project (Mhlambi, 2020; Gyekye, 2004; Sinha and Zulfa, 2023). Inspired by philosophies such as Ubuntu or Ujamaa, DDT should operate in a manner that benefits the society in which they are created and deployed (Eke et al., 2023a; Adelani et al., 2022; Dignum, 2023).
Given the current gap between Africa’s AI readiness and growing interest in AI adoption, many concede external partnership as a necessity (Eke et al., 2023b; African Union, 2024). However, exploitative external relationships set a precedent that curtails African self-determination in data science work (Ndjungu, 2020; Sinha and Zulfa, 2023). When building relationships, there are established obligations that each collaborator owes to the other (Coetzee, 2004b; Metz, 2021). Data collaborations must be predicated on trust, fair attribution of work, and a commitment to prioritizing the agency of African collaborators (Adelani et al., 2022; Nwankwo and Sonna, 2019; Abebe et al., 2021; Gwagwa, 2019; Wareham, 2021; Wiredu, 2004b).
Mend Data-Driven Harms. When disagreements, conflict, or harm occur at any stage of the data science lifecycle, African data ethicists assert the need for mechanisms of accountability and reconciliation to correct wrongs and empower those impacted. In African societies, harm is not just actively making someone’s life worse but also neglecting obligations to the community (Wiredu, 2004a; Gyekye, 2004). A person who causes harm is viewed as a moral failure who must be corrected by their community through sanctions and even mental rehabilitation to address deeper issues connected to their poor actions (Wiredu, 2004b; Coetzee, 2004b). Even the most powerful members of society, such as chiefs, are subject to correction and even dismissal by their community for misconduct (Wiredu, 2004a).
The adoption of AI and other DDT have already caused harm to African populations by way of data bias, socio-economic risk, and privacy violations (Ade-Ibijola and Okonkwo, 2023). There are African data ethicists who stress the need to develop procedures for communities and individuals harmed by DDT to seek restitution (Gwagwa, 2019; Mhlambi and Tiribelli, 2023). These solutions are dependent on African governments and external multinational organizations committing to transparency, equality, and restorative practices (Gwagwa, 2019; African Union, 2024; Okolo et al., 2023; Dignum, 2023; Kiemde and Kora, 2022). African governments can mitigate data harm by being transparent about their potential data collaborations, outlining their plans for data protection before, during, and after the deployment of DDT, and enforcing mechanisms of accountability and dissent from their citizens (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Sinha and Zulfa, 2023). Similar to the dismissal of chiefs, powerful stakeholders acting outside of their agreed duties, must experience restorative consequences, not just a slap on the wrist (Mandaza, 2004; Ndjungu, 2020; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Mhlambi and Tiribelli, 2023; Langat et al., 2020; Gwagwa, 2019; Mhlambi, 2020; Biko, 2004; Coetzee, 2004a).
4.5. Center Communities on the Margins
Community involvement ensures African DDT considers the needs and potential impacts of rural communities, women, youth and populations beyond the end-users.
Design in Solidarity. Solidarity is understood as looking out for other diverse communities based on mutual respect and the goal of social cohesion (Gwagwa, 2019; Mhlambi, 2020). In Ubuntu understanding, solidarity is a deep care for others, including people of the past, present, future, and the environment (Mhlambi and Tiribelli, 2023; Gwagwa, 2019; Okolo et al., 2023; Dignum, 2023; Gwagwa et al., 2022). With this perspective, African data scientists in our corpus call for DDT to not be developed with only the end user in mind but all other communities that could be impacted by the technology (Gwagwa et al., 2022; Olojede, 2023; Gyekye, 2004).
While Africa needs to be included in global data science efforts, Africa itself is home to vastly diverse communities that should also be appropriately represented in these efforts (Adelani et al., 2022; Gwagwa, 2019; Goffi, 2023). African communities’ underrepresentation in datasets across all data science tasks is due to, as Gwagwa described, being “uncounted, unaccounted, and discounted” (Gwagwa, 2019). Leaving communities out of data also excludes them from the benefits DDT provide (Gwagwa et al., 2022). Given the need to build explicitly African DDT, the lack of African datasets is a threat to efficacy (Okolo et al., 2023; Olojede, 2023; Ade-Ibijola and Okonkwo, 2023). Including marginalized communities requires mutual respect for diverse perspectives and creating procedures such as impact assessments to provide opportunities for inclusive input (African Union, 2024; Abebe et al., 2021; Goffi, 2023; Mhlambi, 2020). In addition, it’s important to challenge the social, political, and economic dynamics that push communities to the margins in the first place (Kiemde and Kora, 2022; Olojede, 2023; Segun, 2021; Day and Rennie, 2023; Uzomah et al., 2023; Abebe et al., 2021).
Solidarity violations between African countries is of particular concern. The success of one African community should not be predicated on the suffering of another (Ndjungu, 2020; Biko, 2004). Upholding solidarity means that all actions made in the data science lifecycle should explicitly protect or improve the lives of vulnerable or marginalized communities. Exploiting the vulnerability of another is not only unethical but unsustainable due to our interconnected nature. The suffering of one community will eventually lead to the destruction of all communities (Nwankwo and Sonna, 2019). Banding together, “watching one another’s back,” and developing DDT as a united front is key to mitigating harm (Olojede, 2023; Nyerere, 1962).
Center Remote & Rural Communities. Development, especially technical development, is usually focused in urban centers and excludes remote and rural communities (Ade-Ibijola and Okonkwo, 2023; Sinha and Zulfa, 2023; Osaghae, 2004). Given the lack of infrastructure in remote and rural communities, DDT should be used to develop and optimize infrastructures and public services for these regions (Carman and Rosman, 2023; African Union, 2024). However, it’s important to keep in mind that RDS done on behalf of rural and remote communities that do not consider their culture, livelihoods, and direct input can lead to harm (African Union, 2024; Ndjungu, 2020; Carman and Rosman, 2023).
Center Women. Due to the prevalence of patriarchy in many African societies, authors in our corpus encourage the agency of women in data science efforts. A few documents suggest that women-led technology businesses and the education of women and girls should be incentivized (African Union, 2024). However, open questions remain about how to maintain African women’s participation in a field known to be male-dominated and antagonistic to women (African Union, 2024; Gwagwa, 2019).
Afro-feminists have a response to the techno-chauvinism that dominates data science (Sinha and Zulfa, 2023). Rather than centering women in general, there must be a recognition of the intersectional status of African women (Sinha and Zulfa, 2023). As articulated by Rosebell Kagumire, African women experience domination through systems of patriarchy, race, sexuality, and global imperialism (Dieng, 2023; Coetzee, 2004b). Therefore, data-driven technology should be developed with the complex needs of African women in mind, because their compounded experiences of marginalization provide insight into the needs of various oppressed populations (Sinha and Zulfa, 2023). There are numerous examples of African women harnessing the internet to fill in the gaps of an oppressive society and DDT holds similar potential (Dieng, 2023). For example, Chil AI Lab Group is a women-led data science collective successfully using data technology to address the often neglected health needs of women in Africa (Eke et al., 2023a).